OpenStudy (anonymous):

sqrt x+2 -2=x sqrt is over the x+2 only

OpenStudy (anonymous):

x^2+3x+2

OpenStudy (anonymous):

\[\sqrt{x+2}-2=x\] \[\sqrt{x+2}=x+2\] x+2=(x+2)^2 x^2+4x-x+4-2=x^2+3x+2

OpenStudy (anonymous):

(x+2)(x+1)=0 x=-2 or x=-1